OpenStudy (anonymous):

1,074 kids were asked why they don’t clean their room and 66% percent said it was because they do not like to clean their room. How many kids is this?

OpenStudy (anonymous):

1074* 2/3 = 358*2 = 716

OpenStudy (anonymous):